OpenStudy (anonymous):

Solve And Show Steps 8/X^2-16 Subtracted By 1/X-4 = 1

OpenStudy (whpalmer4):

\[8/(x^2-16) - 1/(x-4) = 1\] \[8/(x^2-16)=8/[(x-4)(x+4)]\] \[8/[(x-4)(x+4)]-(x+4)/[(x-4)(x+4)]=1\] \[(8-(x+4))/[(x-4)(x+4)]=1 = (4-x)/[(x-4)(x+4)] \] multiply both by -1 to convert (4-x) to (x-4) and cancel (x-4) from top and bottom \[-1 = 1/(x+4)\] You should be able to solve it for x from there!